a nurseryvpays 1025 for 11 crates of flowering plants and 13 crates of herbs. later, the nursery pays an additional 450 for 10 more crates of herbs. what is the price of each crate of flowering plants?

Found 2 solutions by Boreal, ikleyn:
Answer by Boreal(15235)

11F+13H=1025
but 10H=450, so H=45
substitute
11F+13(45)=1025
11F=1025-585=440
F=$40 per flowering plants

Answer by ikleyn(52787)

.

From the condition, it should be clear to you that the price per crate of herbs is 45 monetary units.


Therefore, from 1025 monetary units, subtract 13*45 to find the cost of 11 crates of flowering plants.


The last step is to divide this difference by 11 to find the price per crate of flowering plants.

---------------

Having so clear instruction, you MUST be able to complete the assignment on your own.